# Realistic Extreme Image Rescaling via Generative Latent Space Learning

### v1

@@ 3.2.3 Loss Function
Ours 24.39 22.94 0.6515 0.5943 0.3846 0.5058 0.1524 0.2451 58.28 55.68 0.6663 0.6231
3.3 Perception-oriented Optimization
The $\hat{z}$ in the first stage can be passed to the decoder $\mathcal{D}$ to obtain the rescaled HR image. Although this result has high fidelity, it lacks rich texture. Therefore, training in the second stage is designed to enhance perceptual quality. Inspired by relevant content generation [32, 19], we adopt the adapter architecture to guide the generation of the SD model. The adapter module takes the rescaled features $\hat{z}$ as input and extracts multi-scale features in a coarse-to-fine manner, allowing it to output multi-scale features of different sizes:
$\{F_{c}^{1},F_{c}^{2},...,F_{c}^{K}\}=\mathrm{Ada}(\hat{z})$ (13)
Subsequently, these features are element-wisely added to different layers of the denoising U-Net. Considering that image rescaling task places a higher emphasis on fidelity, we employ conditional control over both the encoding and decoding layers of the denoising U-Net:
$F_{\mathrm{enc}}^{i}=F_{\mathrm{enc}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (14)
$F_{\mathrm{dec}}^{i}=F_{\mathrm{dec}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (15)
Here, $F_{\mathrm{enc}}^{i}$ and $F_{\mathrm{dec}}^{i}$ represent the output features of the encoder or decoder of the denoising U-Net at the $i$-th layer. Finally, the SD model and adapter are optimized using the following loss function:
$\mathcal{L_{\mathrm{stage2}}}=\mathrm{\mathbb{E}}_{\left(z,\epsilon,\hat{z},t\right)}\parallel\epsilon-\epsilon_{\theta}\left(z_{t},t,\hat{z}\right)\parallel^{2}$ (16)
Figure 5: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results. Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
Figure 6: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both HR images and corresponding edge maps. Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
@@ 4.1 Experimental Setup
4.1.1 Datasets and Evaluation Metrics
We train the proposed model using the DIV2K dataset [1] and the Flickr2K dataset [17]. Due to the low resolution of conventional benchmark datasets such as Set5, Set14, and BSD100, they are not suitable for extreme rescaling tasks [5]. Therefore, we utilized four high-resolution datasets: DIV2K, Urban100 [6], CLIC2020 [23], and DIV8K [5] to evaluate our model. We adopt PSNR and SSIM [27] for quantitative evaluation. For perceptual quality evaluation, reference-based metrics LPIPS [34] and DISTS [2] are used. For non-reference image quality assessment, we use metrics including MUSIQ [10] and CLIPIQA [24].
@@ 4.1.2 Training Details
Our LSDIR is built based on the SD 2.1-base model and the training is divided into two stages. First, we train the DFRM together with the pre-trained VAE model for 200K iterations using Eq. 12 The learning rate is initialized as $2\times 10^{-4}$ and reduced by half every 50k iterations. After the training of DFRM, we adopt Eq. (16) to train the Adapter and finetune the diffusion U-Net. We only finetune the linear layers of the U-Net for its better performance. During both stages, we adopt Adam optimizer [12] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and set the mini-batch size to 32.
@@ 4.2 Results and Comparison with SOTA
4.2.1 Quantitative Results
Tab. 1 presents the 16$\times$ and 32$\times$ quantitative image rescaling results on four datasets. The performance of SR methods are generally poor. While methods based on INNs achieve the highest PSNR and SSIM scores due to consideration of the correlation between downscaling and upscaling. VQIR exhibits the best LPIPS by fully utilizing the prior provided by pre-trained VQGAN and being trained based on LPIPS-loss. As our proposed method is trained in the latent space, the results exhibit the best semantic integrity, which leads to significant improvements in two non-reference metrics. It is important to note that while our method slightly lags behind VQIR in terms of certain full-reference metrics, it excels in reconstructing accurate semantics and regular structures, as will be demonstrated in the following section.
@@ 4.2.2 Qualitative Results
The quantitative experiments demonstrate that our model exhibits outstanding performance on the Urban100 dataset, confirming that our method produces more accurate structural features. Therefore, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig. 6. Traditional regression models, such as EDSR, RCAN, IRN, and HCFlow, tend to generate overly-smooth results with few sharp edges. Although GAN-based methods, including ESRGAN and VQIR, generate more details, their edge information lacks accuracy and regularity. In contrast, our approach, by leveraging the prior knowledge encapsulated in the Stable Diffusion model, is able to reconstruct more accurate and rich edge details.
@@ 4.3.1 Effectiveness of Feature Rescaling
To verify the effectiveness of the feature rescaling operation, we move the rescaling operation from the latent space to the pixel space. This means we directly train a downsampling network to downscale the HR $x$ to LR $y$. Subsequently, $y$ is used as a condition to guide the generation process of the SD model. As shown the 1st row in Tab. 2, the rescaling operation in the latent space can improve reconstruction performance, as evidenced by lower LPIPS and DISTS.
In Fig. 7, we plot the LPIPS variation curves on the validation set for the two rescaling schemes. The pixel space-based rescaling scheme requires a longer time to converge, whereas the latent space based scheme can stably converge to a better result in a shorter time. Additionally, we observe that the pixel space based scheme experiences severe color shifts in the middle of training, as illustrated in Fig. 8. It requires a substantial number of iterations to learn the global color distribution of the reconstructed images. This is because using pixels as the conditioning input for Stable Diffusion makes it difficult to align with the intrinsic priors of the pre-trained model, leading to issues such as unstable training and longer training time.
Figure 7: The validation LPIPS curve for rescaling models based on latent space and pixel space.
Figure 8: Reconstructed results after 60k training steps. The pixel space based scheme shows severe color shifts.
@@
$\times$ $\checkmark$ $\checkmark$ 0.3818 0.1333
@@
$\checkmark$ $\times$ $\checkmark$ 0.3559 0.1188
4.3.2 Effectiveness of SD Prior
To validate that SD model provides rich priors about real-world HR images, we directly omit the second training stage. As shown in Tab. 2 (2nd row), the introduction of the SD priors can significantly improve performance of LPIPS and DISTS and greatly enhance the perceptual quality of the reconstruction results.
@@ 4.3.3 Effectiveness of DFRM Design
Our DFRM includes three specific designs: the domain converter module, the pixel guidance module, and the stop-gradient operation in the loss function Equ. 11. To analyze their impact, we remove each component individually and plot the validation loss curves, as shown in Fig. 10. It can be observed that the introduction of the INN and the stop-gradient operation can both reduce reconstruction error. Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as shown in Tab. 3. Therefore, the design of our DFRM strikes a good balance between feature reconstruction and LR image quality.
Ideally, the domain converter composed of a set of invertible neural networks should be able to accurately perform bidirectional mapping between the feature domain and the pixel domain, enabling the LR to share the same content with HR images. However, due to the limited representation power of INN, the intermediate features $w$ obtained by simply downscaling the latent code $z$ are not well-suited for conversion to the pixel domain. The introduction of pixel guidance allows for embedding some pixel-domain information into the intermediate features $w$, thereby achieving a balance between reconstruction performance and LR image quality. As shown in Fig. 9, the LR images exhibit significant noise and color distortions in the absence of pixel guidance. However, the introduction of pixel guidance significantly alleviates these issues.
Figure 9: Comparisons of LR images visual quality with and without pixel guidance.
Table 3: Ablation study of pixel guidance module. The introduction of pixel guidance can significantly enhance the quality of LR images.
Method PSNR\ $\uparrow$ SSIM\ $\uparrow$ NIQE\ $\downarrow$
w/o pixel guidance 19.88 0.5898 22.33
with pixel guidance 21.99 0.7237 13.27
4.4 Discussions
4.4.1 Comparison with JPEG Compression
Our model can be combined with existing image compression methods to achieve more extreme compression ratios. To validate this, we use the lossless compression method PNG to compress LR images and compare its performance with JPEG compression at various LR sizes. Specifically, we train rescaling models at three different scales (8$\times$, 16$\times$, and 32$\times$) and record the bpp by the LR images. We also adjust the quality parameters of JPEG compression accordingly for comparison. As shown in Fig. 11, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of JPEG compression. In Fig. 12, we present the visual quality, the bpp and the LPIPS of the reconstructed images. Compared to JPEG compression, our method requires comparable or less storage space while achieving superior objective metrics. Additionally, JPEG compression tends to produce noticeable block artifacts and color distortions in the background regions of the images, whereas our model maintains superior reconstruction quality across all three scales. These comparisons demonstrate the potential of our model for image compression.
@@ 4.4.1 Comparison with JPEG Compression
Figure 12: Visual comparisons between our model and JPEG compression. Below the images, the bpp and LPIPS of the reconstructed images are provided.
4.4.2 Fidelity-realism Trade-off
Our two-stage model outputs two features in the latent space ($\hat{z}$ and $\hat{z}_{0}$), which represent reconstruction results with high fidelity and high realism respectively. It is important to note that the high fidelity mentioned here does not refer to higher pixel fidelity but to semantic fidelity in the latent space. By fusing the results of the two stages, we can achieve a trade-off between fidelity and realism:
$z_{\mathrm{fus.}}=\alpha\hat{z}+\left(1-\alpha\right)\hat{z}_{0}$ (17)
$x_{\mathrm{fus.}}=\mathcal{D}\left(z_{\mathrm{fus.}}\right)$ (18)
As shown in Fig. 13, by adjusting the value of $\alpha$, we can achieve a smooth transition between fidelity and realism. More results are provided in the supplementary materials. As shown in Fig. 14, It can be observed that by adjusting the fusion weights, we can achieve a smooth transition between fidelity and realism.
Figure 13: By adjusting the fusion ratio of two latent codes, we can achieve a trade-off between fidelity and realism.
Figure 14: By adjusting the fusion weights, our model can achieve a trade-off between fidelity and realism. The perceptually optimized results display richer details, but they also exhibit some artifacts.
Figure 15: Visual comparisons of LR images.
@@ 5 Limitations
Although our model demonstrates superior performance in extreme image rescaling tasks, it still has some limitations. One of them is the quality of the LR images. As shown in Fig. 15, the LR images generated by IRN closely resemble the bicubic downsampled image. In contrast, the LR images produced by HCFlow and VQIR exhibit some color shifts. LR images generated by our model are slightly noisy. In the future, more advanced invertible network architectures and more effective loss functions could be explored to further improve the quality of LR images.

### v2

@@ 3.3 One-Step Perception Enhancement
Ours 25.93 24.15 0.7107 0.6302 0.3198 0.4376 0.0927 0.1798 57.75 58.89 0.6598 0.6461
Figure 6: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results. The first row represents the original high-resolution images. Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
Figure 7: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both the rescaled images and corresponding edge maps. Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
Figure 8: Visual comparisons of 32$\times$ rescaling methods on the DIV8K dataset
Figure 9: Visual comparisons of downscaled LR images with sota methods.
In this part, we fine-tune the denoising U-Net network using LoRA, while also jointly training the TPM and refiner module. To achieve a balance between fidelity and realism, we supervise the reconstruction results $\hat{x}$ using pixel loss and perceptual loss as follows:
$\begin{split}\mathcal{L}_{\mathrm{enh.}}=&\underbrace{\left\|\hat{x}-x\right\|_{1}}_{\text{pixel loss}}+\\ &\underbrace{\lambda_{\mathrm{lpips}}\mathcal{L}_{\mathrm{lpips}}\left(\hat{x},x\right)+\lambda_{\mathrm{dists}}\mathcal{L}_{\mathrm{dists}}\left(\hat{x},x\right)}_{\text{perceptual loss}}\end{split}$ (13)
@@ 4.1 Experimental Setup
4.1.1 Datasets and Evaluation Metrics.
We train the proposed OSIRDiff using the DF2K dataset [1, 19]. Considering that traditional benchmarks such as Set5, Set14, and BSD100 often have lower quality and resolution [5], they are not suitable for evaluating extreme rescaling methods. Therefore, we utilize four high-resolution datasets: DIV2K, Urban100 [7], CLIC2020 [32], and DIV8K [5] to evaluate models been compared. We adopt PSNR and SSIM [37] for quantitative evaluation. For perceptual quality evaluation, reference-based metrics LPIPS [49] and DISTS [3] are used. For non-reference image quality assessment, we use metrics including MUSIQ [13] and CLIPIQA [33].
@@ 4.1.2 Training Details
OSIRDiff is built based on the SD 2.1-base model, and the training is divided into three stages. First, we train the DFRM using equation (8). Then we adopt equation (13) to jointly train the LoRA layers, the TPM and the refiner module. Finally, we use a smaller learning rate to jointly fine-tune DFRM, LoRA layers, TPM and refiner module using a combination of equations (8) and (13). More training details are provided in the supplementary material.
@@ 4.2 Results and Comparison with SOTA
We compare the proposed OSIRDiff with three kinds of methods: (1) bicubic downscaling and SR, i.e., EDSR [19], RCAN [50] and ESRGAN [36]; (2) flow-based rescaling models, i.e., IRN [41] and HCFlow [18]; (3) prior-based extreme rescaling method VQIR [39]. Some of them are proposed for 4$\times$ rescaling factors and we extended them to 16$\times$ or 32$\times$. To ensure a fair comparison, all methods are trained until convergence on the DF2K dataset following their original configurations.
4.2.1 Quantitative Results
@@ 4.2.2 Qualitative Results
The quantitative experiments demonstrate that our model exhibits superior performance on the Urban100 [7] dataset, confirming that our method produces more accurate structural features. In this section, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig. 7. Traditional regression models, such as EDSR [19], RCAN [50], IRN [41], and HCFlow [18], tend to generate overly-smooth results. GAN-based methods, including ESRGAN [36] and VQIR [39], generate more details, but their edge information lacks accuracy and regularity. In contrast, our approach, by leveraging the prior knowledge encapsulated in the SD model, is able to reconstruct more accurate and rich edge details. In Fig. 8, we present a visual comparison of 32$\times$ rescaling results on the ultra-high-resolution dataset, DIV8K [5]. It is evident that our method is capable of recovering images with accurate semantics, even at extreme upscaling factors.
To validate the effectiveness of our downscaling scheme, we conduct a visual comparison of the downscaled LR images with state-of-the-art (SOTA) methods, as shown in Fig. 9. Overall, the LR images generated by our method achieve comparable or even superior visual quality to SOTA image rescaling methods. The LR images produced by VQIR exhibit noticeable color shifts. Furthermore, VQIR generated LR images contain significant noise, whereas our method exhibits a certain degree of ringing artifacts. In future work, we plan to further optimize the quality of the LR images.
@@ 4.3.1 Effectiveness of Feature Rescaling
To verify the effectiveness of the feature rescaling operation, we move the rescaling operation from the latent space to the pixel space. Specifically, we use HCFlow [18], the comparative method with the best PSNR performance, to perform the rescaling operation in the pixel domain, followed by perceptual enhancement using the same approach described in Sec. 3.3. As shown in Fig. 11, the rescaling operator in pixel space takes HR image as input and simultaneously outputs both the LR image and the rescaled image. Then, to use SD for perceptual enhancement, we need to map the rescaled image to the latent space using a VAE encoder and perform the denoising process. Finally, the enhanced latent features are mapped back to the pixel space using the VAE decoder. Due to the nonlinear mapping nature of the VAE encoder, the minimum distance in pixel space does not correspond to the minimum distance in latent space. This causes the rescaling operator trained in pixel space to misalign with the prior of the pre-trained SD model, resulting in a degradation of perceptual quality. As shown the 1st row in Table 2, the rescaling operation in the latent space can improve reconstruction performance, as evidenced by much lower LPIPS and DISTS.
Figure 11: Image rescaling within pixel space and latent space with Stable Diffusion (SD) prior.
As shown in Fig. 12, we conduct visual comparisons between rescaling in latent space and pixel space in two different scenarios. Our latent-space rescaling method is able to reconstruct more realistic textures and more accurate structural features. This indicates that latent-space rescaling operation can preserve sufficiently fine-grained contextual information about the HR image, such as structure, texture, and semantics.
Figure 12: Visual comparisons of our method with rescaling in pixel space and without the SD prior.
@@
$\times$ $\checkmark$ $\checkmark$ 0.3630 0.1154
@@
$\checkmark$ $\times$ $\checkmark$ 0.3026 0.0824
4.3.2 Necessity of SD Prior
To validate that SD model provides rich priors about real-world HR images, we directly omit the perceptual enhancement part. As shown in Table 2 (2nd row), the introduction of the SD priors can significantly improve performance of LPIPS and DISTS and greatly enhance the perceptual quality of the reconstruction results. Here, we compare the latent features before and after perceptual enhancement by decoding them to the pixel space using the VAE decoder, as shown in Fig. 12. It can be seen that the latent features before perceptual optimization exhibit noticeable color shifts and blurriness when mapped to the pixel domain. In contrast, the perceptually enhanced latent features display richer textures and more accurate color information. Therefore, leveraging the rich natural image priors in the pre-trained Stable Diffusion model can significantly enhance the visual quality of rescaled images.
@@ 4.3.3 Effectiveness of DFRM Design
Our DFRM includes two specific designs: the domain converter module and the pixel guidance module. To analyze their impact, we remove each component individually and plot the validation loss curves, as shown in Fig. 10. It can be observed that the introduction of the INN can reduce reconstruction error. Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as demonstrated by the lower guidance loss in Fig. 10. As shown in Fig. 13, the LR images exhibit significant noise and color distortions in the absence of pixel guidance. However, the introduction of pixel guidance significantly alleviates these issues. Therefore, the design of our DFRM strikes a good balance between feature reconstruction and LR image quality.
Figure 13: Visual comparisons of LR images with and without pixel guidance.
Figure 14: Ablation study about time alignment.
4.3.4 Effectiveness of Time Alignment
To validate the effectiveness of our proposed time-step alignment strategy, we retrain our model using fixed time-steps. This involves removing the TPM and refiner modules and employing a manually selected time-step to perform a single denoising step based on equation (9). Following the setting of previous work [40], we select fixed time-steps 1 and 999 for comparison. As shown in Fig. 14, when the time-step is set to a smaller value, the model achieves higher fidelity but lower perceptual quality, and the situation is reversed when a larger time-step is chosen. In contrast, our time-step alignment mechanism can dynamically determine the level of generative capability based on the image content, thereby achieving optimal performance in both fidelity and perceptual quality.
5 Discussions
5.1 Comparison with JPEG Compression
Our model can be combined with existing lossless image compression methods to achieve extreme compression ratios. To validate this, we use the lossless compression image format PNG to compress LR images and compare its performance with JPEG compression at various bit-rates. As shown in Fig. 15, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of the JPEG compression. These comparisons demonstrate the potential of our model for image compression.
@@ 5.1 Comparison with JPEG Compression
In Fig. 16, we present the visual quality, the bpp and the LPIPS of the reconstructed images. Compared to JPEG compression, our method requires comparable or less storage space while achieving superior objective metrics. Additionally, JPEG compression tends to produce noticeable block artifacts and color distortions in the background regions of the images, whereas our model maintains superior reconstruction quality across all three scales.
Figure 16: Visual comparisons between our model and JPEG compression. bpp/LPIPS of the reconstructed images are shown below the images.
5.2 Time-step Distribution
To overcome the non-uniform degradation introduced by the rescaling operation, we design a time-step alignment strategy to predict an optimal time-step $t$ based on the degraded latent variable $\hat{z}$, enabling dynamic allocation of generative capability. During the inference phase, our model can perform tiled inference, allowing it to predict different time-steps for various regions of the same image. We discuss the distribution of time-steps at both the single image and dataset scales.
Figure 17: The time-steps $t$ predicted by the TPM on a single image, along with the corresponding latent variable rescaling error map $|z-\hat{z}|$.
First, we visualize the time-steps corresponding to different regions of the image shown in Fig. 17. We also present the latent variable rescaling error map $|z-\hat{z}|$ for comparison. It can be observed that in more complex regions, the rescaling operation introduces larger errors. Correspondingly, these areas also require larger time-steps. This is because complex regions necessitate greater generative capability for restoration.
Figure 18: The distribution of time-steps predicted by the TPM on the DIV2K validation set at different rescaling scale factors. Smaller scale factor results in smaller time-steps.
Next, we present the distribution of predicted time-steps at different rescaling scales on the DIV2K validation set shown in Fig. 18. The vertical axis and the size of the bubbles represent the mean and variance of the time-steps, respectively. It can be observed that the mean of the predicted time-steps increases as the rescaling scale becomes larger. This is because larger rescaling scales result in greater image degradation, necessitating stronger generative capability for restoration. Additionally, we find that the predicted time-steps for the 16$\times$ rescaling exhibit the largest variance. This indicates that 16$\times$ rescaling is a task that balances both fidelity and generative capacity. In contrast, 8$\times$ rescaling tends to prioritize fidelity, while 32$\times$ rescaling leans more towards generative capabilities.
@@ 6 Limitations
Table 3: The performance comparison between our 4×4$\times$4 × rescaling model and VAE.
Method PSNR $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
VAE 29.00 0.1570 65.54 0.6234
Ours-4×4$\times$4 × 29.32 0.1456 65.47 0.6313
6.1 Upper Bound of OSIRDiff
Since our method performs the rescaling operation in the latent space of a pre-trained VAE, its performance ceiling is inherently constrained by the VAE itself. The VAE encoding is a lossy process and can be described as follows:
$z=\mathcal{E}(x)$ (14)
Here, $x\in R^{H\times W\times 3}$ represents original HR image and $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ represents corresponding latent features. We can measure the compression rate from the perspective of the number of bits:
[…]
$=\left(H\times W\times 3\right)\times 8$ (15)
[…]
$=\left(\frac{H}{8}\times\frac{W}{8}\times 4\right)\times 16$ (16)
$\mathrm{Rate}$ $=\frac{\mathrm{Bits(z)}}{\mathrm{Bits(x)}}=\frac{1}{24}$ (17)
It can be observed that VAE encoding can be regarded as a compression process with a compression ratio of 24. From the perspective of rescaling, this is equivalent to a rescaling operation with a scale factor of $S=\sqrt{24}$. Any rescaling scale smaller than $S$ will inevitably be limited by the reconstruction performance of VAE, as the VAE encoding process has already discarded significant amounts of information.
Furthermore, we directly encode and decode input HR images using VAE without performing any rescaling operations, representing the upper performance bound of our model. Additionally, we train an OSIRDiff model for 4×4$\times$4 × rescaling and compare its performance with the reconstruction capability of the VAE, as illustrated in Table 3. It can be observed that the performance of our 4×4$\times$4 × rescaling model is very close to the reconstruction performance of the VAE, indicating that the rescaling performance at this scale is limited by the VAE reconstruction capacity. This also highlights a potential limitation of SD prior itself that it only demonstrates significant advantages when handling highly ill-posed tasks [8, 2, 46]. Consequently, our proposed OSIRDiff is similarly affected by this limitation and performs well primarily at larger rescaling scales (>$S$). In the future, we could explore the incorporation of shallow features into the VAE decoding process to enhance its fidelity, thereby enabling the OSIRDiff framework to be more applicable to small-scale rescaling tasks.
Table 4: Comparison of inference time and parameter sizes across methods.
Time (s) Total Parameters (M) Trainable Parameters (M)
IRN 0.359 34.2 34.2
HCFlow 0.363 7.24 7.24
VQIR 0.897 92.9 20.7
Ours 0.711 1352.5 63.2
6.2 Inference Time and Model Size
In Table 4, we compare the inference speed and parameter sizes of different methods. All tests are performed on an RTX 3090 using images with a resolution of $1024\times 1024$. Due to leveraging the SD architecture, our method has the largest total number of parameters and the highest number of trainable parameters. However, since our approach only requires a single denoising step, its inference time does not increase significantly compared to other methods. The inference speed of our method even surpasses that of the VQGAN-based [4] method, i.e. the VQIR. This is because, compared to VAE, the latent space of VQGAN has a higher number of channels. Additionally, the nearest-neighbor search process in VQGAN incurs significant computational overhead, adversely affecting its inference speed. In the future, we plan to explore architecture distillation techniques to further reduce the size of our image rescaling model and improve its inference speed.